One tower is 200 ft. high and is 250 feet away from another building that is 150 ft. high. At the same time, two crows, one from each tower, fly off at the same speed heading toward food that is on a level, straight road located between the two towers. The crows reach the food at the same time. How far is the grain from the foot of the first tower (the 200ft one)?
Answer by scott8148(6628)

let x=distance

200^2+x^2=150^2+(250-x)^2 __ 200^2+x^2=150^2+250^2-500x+x^2

200^2-150^2-250^2=-500x __ 40000-22500-62500=500x __ -45000=-500x __ 90=x